# A Systematic Survey on Large Language Models for Algorithm Design

### v2

@@ Abstract.
Algorithm Design (AD) is crucial for effective problem-solving across various domains. The advent of Large Language Models (LLMs) has notably enhanced the automation and innovation within this field, offering new perspectives and promising solutions. Over the past three years, the integration of LLMs into AD (LLM4AD) has progressed significantly, finding applications in diverse areas such as optimization, machine learning, mathematical reasoning, and scientific discovery. Given the rapid development and broadening scope of this field, a systematic review is both timely and essential. This paper provides a systematic review of the works on LLM4AD. First, we present an overview and summary of existing studies. Then, we present a systematic taxonomy and a review of existing works along four dimensions, including the role of LLMs, search techniques, prompt strategies, and applications, with a discussion of the potential and achievements of using LLMs. Finally, we explore current challenges and propose several open questions and promising directions for future research.
Algorithm design, Large language model, Optimization, Heuristic, Hyperheuristic, Evolutionary algorithm.
@@ 1. Introduction
Algorithms play a crucial role in addressing various problems across diverse fields, including industry, economics, healthcare, and technology (Kleinberg, 2006; Cormen et al., 2022). The traditional hand-crafted approach to algorithm design requires extensive expert knowledge and considerable effort, which is tedious and time-consuming. There has been a growing interest in the integration of learning and computational intelligence techniques to streamline and enhance the algorithm development process (Bengio et al., 2021; Tang and Yao, 2024).
@@
In recent years, Large Language Models (LLMs) have emerged as a significant breakthrough in generative artificial intelligence. LLMs are distinguished by their huge model size, extensive training data, and remarkable performance in various research areas, including mathematical reasoning (Ahn et al., 2024), code generation (Jiang et al., 2024), and scientific discovery (Wang et al., 2023a).
@@
In the past three years, the integration of Large Language Models for Algorithm Design (LLM4AD) has become an emerging area of research that promises to revolutionize how algorithms are conceived, optimized, and implemented. The impressive power and adaptability of LLMs have shown promising in improving and transforming the algorithm design process, such as heuristic generation (Liu et al., 2024d), code optimization (Ma et al., 2024a), and even the creation of novel algorithms (Girotra et al., 2023) tailored to specific problems. This approach not only reduces the human effort required in the design phase but also enhances the creativity and efficiency of the produced solutions (Liu et al., 2024d; Romera-Paredes et al., 2024).
@@
While LLM4AD is gaining traction, systematic reviews in this nascent field remain sparse. Existing literature primarily focuses on applications of LLMs in specific algorithmic contexts. For example, several papers have been carried out to survey the LLM for/with specific optimization topic (Huang et al., 2024b; Wu et al., 2024a; He et al., [n. d.]) and others review the general LLM applications (Hadi et al., 2024) or the application on specific problems including electronic design automation (Zhong et al., 2023), planning (Pallagani et al., 2024), soft engineering (Hou et al., 2023), recommendation (Wu et al., 2024b) and agent (Wang et al., 2024c). This paper aims to fill this gap by providing a systematic review with a multi-dimensional taxonomy of the current state of LLMs in algorithm design. We will explore various applications, discuss key challenges, and propose directions for future research. By synthesizing these insights, this paper contributes to a deeper understanding of the potential of LLMs in enhancing algorithm design and lays the groundwork for further innovations in this exciting field. We expect it to be a helpful resource for both new researchers entering the field and experienced experts looking for a consolidated and systematic update on current developments. The contributions of this paper are as follows:
@@
  • Discussion on Challenges and Future Directions: We go beyond mere summarization of existing literature to critically analyze the limitations present in current research on LLMs for algorithm design. Furthermore, we propose potential future research directions, including developing domain-specific LLMs, exploring multi-modal LLMs, facilitating human-LLM interaction, using LLMs for algorithm assessment and understanding LLM behavior, advancing fully automatic algorithm design, and benchmarking for systematic evaluation of LLMs in algorithm design. This discussion is intended to spur novel approaches and foster further advancements in the field.
@@ 2.1. Scope of Survey
This paper aims to conduct a systematic survey and classification of existing research works in the emerging field of Large Language Model for Algorithm Design (LLM4AD). We do not intend to cover all the literature on LLMs and algorithms. We delineate the scope of our survey as follows:
  • The term Large Language Models refers to language models of sufficient scale. These models typically utilize a transformer architecture and operate in an autoregressive manner (Zhao et al., 2023). 1) Studies employing smaller models for algorithm design, such as those in conventional model-based and machine learning-assisted algorithms (Bengio et al., 2021), are excluded. While it is challenging to precisely define “large” models, most leading-edge LLMs contain over one billion parameters (Zhao et al., 2023; Minaee et al., 2024). 2) Research utilizing other large models that lack language processing capabilities, such as purely vision-based models, are not considered. However, multi-modal LLMs that include language processing are within our scope.
  • The term Algorithm in this context refers to a set of mathematical instructions or rules designed to solve a problem, particularly when executed by a computer (Cormen et al., 2022). This broad definition encompasses traditional mathematical algorithms (Ahn et al., 2024), most heuristic approaches (Mart et al., 2018; Michel Gendreau, 2019), and certain agents or policies that can be interpreted as algorithms (Xi et al., 2023).
@@
[…]
Four stages for paper collection.
Figure 1. Four stages for paper collection.
@@ 2.3. Taxonomy
This paper presents a taxonomy organized into four dimensions: 1) LLM Roles, 2) Search Methods, 3) Prompt Methods, and 4) Applications.
@@ 3.1. LLMs as Optimizers (LLMaO)
In LLMaO, LLMs are utilized as a black-box optimizer within an algorithm framework to generate and refine solutions (Fig. 4). The integration of LLMs into optimization tasks leverages their ability to understand and generate complex patterns and solutions and good flexibility in prompt engineering (Yang et al., 2024; Liu et al., 2023a, 2024c), which might be infeasible for traditional optimizers. However, due to their black-box nature, they usually lack interpretability and face challenges in large-scale generation (Liu et al., 2023a).
@@
Differing from direct solution generation, Lange et al. (2024) investigate the use of LLMs in designing evolution strategies, introducing a new prompting strategy to enhance the mean statistic in their EvoLLM method, which shows superior performance over baseline algorithms in synthetic black-box optimization functions and neuroevolution tasks. They also demonstrate that fine-tuning LLMs with data from teacher algorithms could further improve the performance of EvoLLM. Moreover, Custode et al. (2024) present a preliminary study that uses LLMs to automate hyperparameter selection by analyzing optimization logs and providing real-time recommendations.
Beyond traditional optimization tasks, LLMaO has been widely adopted in prompt engineering for LLMs, a process often referred to as “automatic prompt optimization” (Zhou et al., 2022). These methods primarily involve iterative refinement of prompts by LLMs to improve their effectiveness for specific models (typically LLMs). Techniques include resampling-based strategies, where LLMs generate variations of original prompts while maintaining semantic similarity (Wang et al., 2023b), and reflection-based strategies, where LLMs optimize by analyzing and learning from previous prompt iterations or errors (Guo et al., 2023), have been explored. Ma et al. (2024d) note that LLM optimizers often struggle to accurately identify the root causes of errors during the reflection process, influenced by their pre-existing knowledge rather than an objective analysis of mistakes. To address these issues, they proposed a new approach termed “automatic behavior optimization”, aimed at directly and more effectively controlling the behavior of target models.
@@ 3.2. LLMs as Predictors (LLMaP)
LLMaP employs LLMs as surrogate models (Fig. 5) to predict the outcomes or responses of solutions, functioning either in a classification or regression context (Hao et al., 2024). Compared to other model-based predictors, such as the Gaussian process and conventional neural networks, 1) LLMs are capable of processing and generating human-like responses based on the training they have received on vast datasets. This capability allows them to understand and interpret complex patterns in data, making them suitable for tasks where traditional modeling techniques might struggle due to the intricacies in the data and complicated representations (Egami et al., 2024). 2) Pre-trained LLMs can significantly reduce the computational load and time required compared to training high-fidelity models (Jawahar et al., 2023; Hao et al., 2024).
@@
The majority of LLMaP works use LLMs as pre-trained models to predict solution scores. For instance, LLMs have been used as performance predictors for deep neural network architectures by Jawahar et al. (2023). It offers a cost-effective alternative for performance estimation in neural architecture search. Zhang et al. (2024e) introduce LINVIT, an algorithm that incorporates guidance from LLMs as a regularization factor in value-based RL to improve sample efficiency. Science discovery is another domain that LLMaP has commonly investigated. For example, Li et al. (2023) introduce CodonBERT for sequence optimization of mRNA-based vaccines and therapeutics. CodonBERT uses codons as inputs and was trained on over 10 million mRNA sequences from various organisms. Soares et al. (2024) demonstrate the use of transformer-based models in predicting the performance of battery electrolytes. Other applications include employing LLMs to determine the fame score of celebrities to predict the box office performance of projects in the motion pictures industry (Alipour-Vaezi and Tsui, 2024) and adopting LLMs to score the video question answering by using detailed video captions as content proxies (Zhang et al., 2024a).
@@
For classification, Hao et al. (2024) introduce LAEA, which employs LLMs as surrogate models within evolutionary algorithms for both regression and classification, eliminating the need for costly model training. In another study, Chen et al. (2023b) develope a label-free node classification method that leverages LLMs to annotate nodes. These annotations are subsequently used to train graph neural networks, resulting in enhanced performance. Moving beyond binary classification, Bhambri et al. (2024) utilize LLMs to predict discrete actions for constructing reward shaping functions in Reinforcement Learning (RL). Their method demonstrated effectiveness within the BabyAI environment, showcasing the versatility of LLMs in various settings. Wang et al. (2024d) explore the use of LLMs in federated search, applying them in a zero-shot setting to effectively select resources. This approach highlights the potential of LLMs in improving resource selection without prior explicit training on specific tasks. Lastly, Mehrdad et al. (2024) focus on automating the relevance judgment of query-item pairs in product searches. By finetuning LLMs, they achieved significant improvements, underscoring the adaptability and effectiveness of LLMs in complex search tasks.
LLMs can respond to more than numerical data, which are quite different from conventional prediction models. For example, Egami et al. (2024) utilize LLMs to generate imperfect annotation surrogates for documents in computational social science research, combined with high-quality labels to provide valid statistical inference for downstream analyses. In addition, the multi-modal LLMs demonstrate exceptional capability in utilizing multi-model information. An evolutionary strategy with a vision-language model has been developed by Wong et al. (2024b) for practical 3D car designs. The vision LLM is used to predict practicality degree for designs that satisfy the given text prompt.
@@ 3.3. LLMs as Extractors (LLMaE)
LLMaE employs LLMs to mine and extract embedded features or specific knowledge from target problems and/or algorithms, which are then utilized in the enhancement of algorithm-based problem solving (Fig 6). This process leverages the unique capabilities of LLMs, such as text and code comprehension, allowing them to discern subtle patterns and relationships within the data that might not be evident through conventional feature extraction methods (Wu et al., 2024c).
@@
In the context of feature extraction, Kristiadi et al. (2024) explore the use of LLMs in accelerating Bayesian optimization in the molecular space. LLMs are used as fixed feature extractors for standard Bayesian optimization surrogate models, and parameter-efficient finetuning methods are leveraged to obtain the posterior of the LLM surrogate.
Beyond conventional feature extraction, LLMs excel in text comprehension and knowledge extraction. An example in LLM-assisted algorithm selection by Wu et al. (2024c) utilize LLMs to extract high-dimensional algorithm representations by comprehending code text. These representations are combined with problem representations to determine the most suitable algorithm for a specific problem. In another innovative application, Du et al. (2024c) propose a mixture-of-experts framework augmented with LLMs to optimize various wireless user tasks. The LLM is used to analyze user objectives and constraints, thus selecting specialized experts, and weighing decisions from the experts, reducing the need for training new models for each unique optimization problem. Additionally, Becker et al. (2024) discuss the use of LLMs in the automotive and supplier industries, specifically focusing on retrieval-augmented generation systems to improve information retrieval from technical documentation. Memduhoğlu et al. (2024) use LLM embeddings to enhance the classification of urban building functions by interpreting OpenStreetMap tags and integrating them with physical and spatial metrics. Traditional techniques, which previously struggled with semantic ambiguities, are outperformed by LLMs due to their superior ability to capture broader language contexts.
@@ 3.4. LLMs as Designers (LLMaD)
LLMaD directly creates algorithms or specific components (Fig. 7). This utilization of LLMs extends their application beyond traditional boundaries, enabling them to actively participate in algorithm development by generating heuristics (Liu et al., 2024d), writing code snippets (Ma et al., 2024b), or formulating functions (Romera-Paredes et al., 2024) that can be integrated into algorithmic systems. By doing so, LLMs can significantly accelerate the algorithm design process, reduce human effort, and bring creativity and optimization to algorithm development (Liu et al., 2024d), which is difficult to achieve through conventional methods alone.
@@
Function design is among the early applications of LLMaD. Eureka (Ma et al., 2024b) leverages the capabilities of LLMs in code-writing, and in-context learning to evolve and optimize reward functions for RL. It can generate reward functions without specific prompts or predefined templates, achieving better performance than rewards designed by human experts. Similarly, Auto MC-Reward (Li et al., 2024b) utilizes LLMs to automatically design dense reward functions for RL agents in environments with sparse rewards. The three key components of Auto MC-Reward work together to iteratively refine the reward function based on feedback from the agent’s interactions with the environment. Through this iterative process, the agent is able to learn complex tasks more efficiently, as demonstrated in experiments in Minecraft. Moreover, FunSearch (Romera-Paredes et al., 2024) adopts LLMs for function generation in an evolutionary framework with a multi-iland population management. FunSearch demonstrated promising results on both mathematical problems and combinatorial optimization problems.
@@
EoH (Liu et al., 2024d), originates from AEL (Liu et al., 2023b), presents an early attempt to adopt LLM as a designer for automatic heuristic design. It uses both heuristic ideas and code implementations to represent heuristics and adopts LLM in an evolutionary framework to create, combine, and revise the algorithmic heuristic. EoH is applied to the key heuristic component design for combinatorial optimization problems, where it surpasses both traditional human-designed metaheuristics and deep-learning-based neural solvers. The application of EoH has expanded to include cost-aware acquisition function design for Bayesian optimization (Yao et al., 2024a), adversary attack algorithm design for image adversary attack (Guo et al., 2024b), and heuristic design for edge server task scheduling Yatong et al. (2024). LLaMEA (van Stein and Bäck, 2024) develops an iterative framework to generate, mutate, and select algorithms based on performance metrics and runtime evaluations. The automatically designed algorithms outperform state-of-the-art optimization algorithms on some benchmark instances. ReEvo (Ye et al., 2024a) introduces an evolutionary framework with both short and long-term reflections, which provides a search direction to explore the heuristic space. With the reflection, better results on black-box cases (no problem-specific information is provided in the prompts) have been observed. Moreover, Yan et al. (2024) explore the use of LLMs to generate drafts of high-quality algorithms for electronic health records phenotyping, aiming to improve efficiency and scalability in observational health research and Hazra et al. (2024) combine LLMs with heuristic planning to generate cost-effective plans, showing superior performance in LLM planning approaches. Unlike previous studies that focus on optimizing a single performance criterion, MEoH (Yao et al., 2024b) considers multiple performance metrics, including optimality and efficiency, and seeks a set of trade-off algorithms in a single run in a multi-objective evolutionary framework. A dominance-dissimilarity score is designed for effectively searching the complex algorithm space.
@@
LLM-based agent design has also gained much attention. For example, ADAS (Hu et al., 2024) proposes an automated design of agentic systems, which aims to automatically generate powerful agentic system designs by using meta agents that program new agents. They present a novel algorithm, meta agent search, which iteratively creates new agents from an archive of previous designs, demonstrating through experiments that these agents can outperform state-of-the-art hand-designed agents across various domains. Further studies on LLM-based agent systems are discussed in (Sun et al., 2024a) and (Liu et al., 2024e).

### v3

@@ Abstract.
Algorithm Design (AD) is crucial for effective problem-solving across various domains. The advent of Large Language Models (LLMs) has notably enhanced the automation and innovation within this field, offering new perspectives and promising solutions. Over the past three years, the integration of LLMs into AD (LLM4AD) has seen substantial progress, with applications spanning optimization, machine learning, mathematical reasoning, and scientific discovery. Given the rapid advancements and expanding scope of this field, a systematic review is both timely and necessary. This paper provides a systematic review of LLM4AD. First, we offer an overview and summary of existing studies. Then, we introduce a taxonomy and review the literature across four dimensions: the roles of LLMs, search methods, prompt methods, and application domains with a discussion of potential and achievements of LLMs in AD. Finally, we identify current challenges and highlight several promising directions for future research.
Large language model, Automated algorithm design, Optimization, Heuristic, Hyperheuristic, Evolutionary algorithm.
@@ 1. Introduction
Algorithms play a crucial role in addressing various problems across various domains such as industry, economics, healthcare, and technology (Kleinberg, 2006; Cormen et al., 2022). Traditionally, designing algorithm has been a labor-intensive process that demands deep expertise. Recently, there has been a surge in interest towards employing learning and computational intelligence methods techniques to enhance and automate the algorithm development process (Bengio et al., 2021; Tang and Yao, 2024).
@@
In the realm of artificial intelligence, Large Language Models (LLMs) have marked a significant advancement. Characterized by their vast scale, extensive training, and superior performance, LLMs have made notable impacts in the fields such as mathematical reasoning (Ahn et al., 2024), code generation (Jiang et al., 2024), and scientific discovery (Wang et al., 2023a).
@@
Over the past three years, the application of Large Language Models for Algorithm Design (LLM4AD) has merged as a promising research area with the potential to fundamentally transform the ways in which algorithms are designed, optimized and implemented. The remarkable capability and flexibility of LLMs have demonstrated potential in enhancing the algorithm design process, including performance prediction (Hao et al., 2024), heuristic generation (Liu et al., 2024e), code optimization (Hemberg et al., 2024), and even the invention of new algorithmic ideas (Girotra et al., 2023) specifically tailored to target tasks. This approach not only reduces the human effort required in the design phase but also enhances the creativity and efficiency of the produced solutions (Liu et al., 2024e; Romera-Paredes et al., 2024).
@@
While LLM4AD is gaining traction, there is a notable absence of a systematic review in this emerging field. The existing literature primarily focuses on the applications of LLMs within specific algorithmic contexts. For instance, several studies have been conducted to survey the use of LLMs for optimization topics (Huang et al., 2024b; Wu et al., 2024a; He et al., [n. d.]), while others review general LLM applications (Hadi et al., 2024) or their use in particular domains such as electronic design automation (Zhong et al., 2023), planning (Pallagani et al., 2024), recommendation systems (Wu et al., 2024b), and agents (Wang et al., 2024c). This paper aims to address this gap by providing a systematic review with a multi-dimensional taxonomy of the current state of LLMs in algorithm design. We will also explore various applications, discuss key challenges, and propose directions for future research. By synthesizing these insights, this paper contributes to a deeper understanding of the potential of LLMs to enhance and automate algorithm design and lays the groundwork for further innovations in this exciting field. We expect this paper to be a helpful resource for both newcomers to the field and experienced experts seeking a consolidated and systematic update on current developments. The contributions of this paper are outlined as follows:
@@
  • Discussion on Challenges and Future Directions: We go beyond mere summarization of existing literature to critically analyze the limitations present in current research on LLMs for algorithm design. Furthermore, we highlight potential future research directions, including developing domain-specific LLMs, exploring multi-modal LLMs, facilitating human-LLM interaction, using LLMs for algorithm assessment and understanding LLM behavior, advancing fully automated algorithm design, and benchmarking for systematic evaluation of LLMs in algorithm design. This discussion is intended to spur novel approaches and foster further advancements in the field.
@@ 2.1. Scope of Survey
This paper aims to conduct a systematic survey and classification of existing research works in the emerging field of Large Language Model for Algorithm Design (LLM4AD). We do not intend to cover all the literature on both LLMs and algorithms. We delineate the scope of our survey as follows:
  • The term Large Language Models refers to language models of sufficient scale. These models typically utilize a transformer architecture and operate in an autoregressive manner (Zhao et al., 2023). Studies employing smaller models for algorithm design, such as conventional model-based and machine learning-assisted algorithms (Bengio et al., 2021), are excluded. Research utilizing other large models that lack language processing capabilities, such as purely vision-based models, are not considered. However, multi-modal LLMs that include language processing are within our scope.
  • The term Algorithm in this context refers to a set of mathematical instructions or rules designed to solve a problem, particularly when executed by a computer (Cormen et al., 2022). This broad definition encompasses traditional mathematical algorithms (Ahn et al., 2024), most heuristic approaches (Michel Gendreau, 2019), and certain agents or policies that can be interpreted as algorithms (Xi et al., 2023).
@@
[…]
Four stages for paper collection.
Figure 1. Four stages for paper collection.
2.2. Overview
@@ 2.3. Taxonomy
This paper presents a taxonomy organized into four dimensions as shown in Fig. 4: 1) LLM Roles, 2) Search Methods, 3) Prompt Techniques, and 4) Applications.
@@ 3. LLM Roles
According to the roles of LLM in algorithm design, existing works can be categorized into four classes: LLMs as Optimizers (LLMaO), LLMs as Predictors (LLMaP), LLMs as Extractors (LLMaE), and LLMs as Designers (LLMaD). This section presents the progress and explore the advantages and limitations associated with each category.
@@ 3.1. LLMs as Optimizers (LLMaO)
In LLMaO, LLMs are utilized as a black-box optimizer within an algorithm framework to generate and refine solutions (Fig. 5). The integration of LLMs into optimization tasks leverages their ability to understand and generate complex patterns and solutions and good flexibility (Yang et al., 2024; Liu et al., 2023a, 2024c). These capabilities are often challenging for traditional optimizers and models to match. However, the black-box nature of LLMs typically results in a lack of interpretability and presents difficulties in large-scale generation (Liu et al., 2023a).
@@
Differing from direct solution generation, Lange et al. (2024) investigate the use of LLMs in designing evolution strategies, introducing a new prompting strategy to enhance the mean statistic in their EvoLLM method, which shows superior performance over baseline algorithms in synthetic black-box optimization functions and neuroevolution tasks. They also demonstrate that fine-tuning LLMs with data from teacher algorithms can further improve the performance of EvoLLM. Moreover, Custode et al. (2024) present a preliminary study that uses LLMs to automate hyperparameter selection by analyzing optimization logs and providing real-time recommendations.
Beyond traditional optimization tasks, LLMaO has been widely adopted in prompt engineering for LLMs, a process often referred to as “automatic prompt optimization” (Zhou et al., 2022). These methods primarily involve iterative refinement of prompts by LLMs to improve their effectiveness for specific models (typically LLMs). Techniques include resampling-based strategies, where LLMs generate variations of original prompts while maintaining semantic similarity (Wang et al., 2023b), and reflection-based strategies, where LLMs optimize by analyzing and learning from previous prompt iterations or errors (Guo et al., 2023), have been explored. Ma et al. (2024d) note that LLM optimizers often struggle to accurately identify the root causes of errors during the reflection process, influenced by their pre-existing knowledge rather than an objective analysis of mistakes. To address these issues, they propose a new approach termed “automatic behavior optimization”, aimed at directly and more effectively controlling the behavior of target models.
@@ 3.2. LLMs as Predictors (LLMaP)
LLMaP employs LLMs as surrogate models (Fig. 6) to predict the outcomes or responses of solutions, functioning either in a classification or regression context (Hao et al., 2024). Compared to other model-based predictors, such as the Gaussian process and conventional neural networks, 1) LLMs are capable of processing and generating human-like responses. This capability allows them to understand and interpret complex patterns in data, making them suitable for tasks where traditional modeling techniques might struggle due to the intricacies in the data and complicated representations (Egami et al., 2024; Wong et al., 2024b). 2) Pre-trained LLMs can significantly reduce the computational load and time required compared to training high-fidelity models (Jawahar et al., 2023; Hao et al., 2024).
@@
The majority of LLMaP works use LLMs as pre-trained models to predict solution scores. For instance, LLMs have been used as performance predictors for deep neural network architectures by Jawahar et al. (2023). It offers a cost-effective alternative for performance estimation in neural architecture search. Zhang et al. (2024f) introduce LINVIT, an algorithm that incorporates guidance from LLMs as a regularization factor in value-based RL to improve sample efficiency. Science discovery is another domain that LLMaP has commonly investigated. For example, Li et al. (2023) introduce CodonBERT for sequence optimization of mRNA-based vaccines and therapeutics. CodonBERT uses codons as inputs and is trained on over 10 million mRNA sequences from various organisms. Soares et al. (2024) demonstrate the use of LLMs in predicting the performance of battery electrolytes. Other applications include employing LLMs to determine the fame score of celebrities to predict the box office performance of projects in the motion pictures industry (Alipour-Vaezi and Tsui, 2024) and adopting LLMs to score the video question answering by using detailed video captions as content proxies (Zhang et al., 2024a).
@@
For classification, Hao et al. (2024) introduce LAEA, which employs LLMs as surrogate models within evolutionary algorithms for both regression and classification, eliminating the need for costly model training. In another study, Chen et al. (2024b) develope a label-free node classification method that leverages LLMs to annotate nodes. These annotations are subsequently used to train graph neural networks, resulting in enhanced performance. Moving beyond binary classification, Bhambri et al. (2024) utilize LLMs to predict discrete actions for constructing reward shaping functions in Reinforcement Learning (RL). Their method demonstrate effectiveness within the BabyAI environment, showcasing the versatility of LLMs in various settings. Wang et al. (2024d) explore the use of LLMs in federated search, applying them in a zero-shot setting to effectively select resources. This approach highlights the potential of LLMs in improving resource selection without prior explicit training on specific tasks. Lastly, Mehrdad et al. (2024) focus on automating the relevance judgment of query-item pairs in product searches. By finetuning LLMs, they have achieved significant improvements, underscoring the adaptability and effectiveness of LLMs in complex search tasks.
Despite these advantages, using LLMs as predictors for algorithm design also suffers from a lack of interpretability, and the results can be influenced by the quality of prompt engineering (Chen et al., 2024b). Moreover, when the target landscape is easily understood, conventional machine learning techniques and surrogate models are often preferred. In such scenarios, LLMs can enhance the utilization of existing modeling methods, for example, by aiding in the selection of the most effective model (Rios et al., 2024).
@@ 3.3. LLMs as Extractors (LLMaE)
LLMaE employs LLMs to mine and extract embedded features or specific knowledge from target problems and/or algorithms, which are then utilized in the enhancement of algorithm-based problem solving (Fig 7). For example, Kristiadi et al. (2024) use LLMs as pre-trained feature extractors and the embeddings are used to enhance standard Bayesian optimization surrogate models in the molecular space.
@@
Beyond embedding-based feature extraction, LLMs excel in text comprehension and knowledge extraction, allowing them to discern subtle patterns and relationships within the data that might not be evident through conventional feature extraction methods. For example, Wu et al. (2024c) utilize LLMs to extract high-dimensional algorithm representations by comprehending code text. These representations are combined with problem representations to determine the most suitable algorithm for a specific problem. Du et al. (2024c) propose a mixture-of-experts framework augmented with LLMs to optimize various wireless user tasks. The LLM is used to analyze user objectives and constraints, thus selecting specialized experts, and weighing decisions from the experts, reducing the need for training new models for each unique optimization problem. Additionally, Becker et al. (2024) discuss the use of LLMs in the automotive and supplier industries, specifically focusing on retrieval-augmented generation systems to improve information retrieval from technical documentation. Memduhoğlu et al. (2024) use LLM to enhance the classification of urban building functions by interpreting OpenStreetMap tags and integrating them with physical and spatial metrics. Traditional techniques, which have previously struggled with semantic ambiguities, are outperformed by LLMs due to their superior ability to capture broader language contexts.
@@ 3.4. LLMs as Designers (LLMaD)
LLMaD directly creates algorithms or specific components (Fig. 8). This utilization of LLMs extends their application beyond traditional boundaries, enabling them to actively participate in algorithm development by generating heuristics (Liu et al., 2024e), writing code snippets (Hemberg et al., 2024), or formulating functions (Romera-Paredes et al., 2024) that can be integrated into algorithmic systems. By doing so, LLMs can significantly accelerate the algorithm design process, reduce human effort, and bring creativity and optimization to algorithm development (Liu et al., 2024e), which is difficult to achieve through traditional methods.
@@
Function design is among the early applications of LLMaD. Eureka (Ma et al., 2024b) leverages the capabilities of LLMs in code-writing, and in-context learning to evolve and optimize reward functions for RL. It can generate reward functions without specific prompts or predefined templates, achieving better performance than rewards designed by human experts. Similarly, Auto MC-Reward (Li et al., 2024b) utilizes LLMs to automatically design dense reward functions for RL agents in environments with sparse rewards. The three key components of Auto MC-Reward work together to iteratively refine the reward function based on feedback from the agent’s interactions with the environment. Through this iterative process, the agent is able to learn complex tasks more efficiently, as demonstrated in experiments in Minecraft. Moreover, FunSearch (Romera-Paredes et al., 2024) adopts LLMs for function generation in an evolutionary framework with a multi-island population management. It demonstrates promising results on both mathematical problems and combinatorial optimization problems.
@@
EoH (Liu et al., 2024e), originates from AEL (Liu et al., 2023b), presents an early attempt to adopt the LLM as a designer for automated heuristic design. It uses both heuristic ideas and code implementations to represent heuristics and adopts LLM in an evolutionary framework to create, combine, and revise the heuristics. EoH is applied on well-studied combinatorial optimization problems, where it surpasses both traditional human-designed metaheuristics and deep-learning-based neural solvers. The application of EoH has expanded to Bayesian optimization (Yao et al., 2024a), image adversary attack (Guo et al., 2024b), and edge server task scheduling (Yatong et al., 2024), among others. Moreover, LLaMEA (van Stein and Bäck, 2024) develops an iterative framework to generate, mutate, and select algorithms based on performance metrics and runtime evaluations. The automatically designed algorithms outperform state-of-the-art optimization algorithms on some benchmark instances. ReEvo (Ye et al., 2024a) introduces an evolutionary framework with both short and long-term reflections, which provides a search direction to explore the heuristic space. With the reflection, better results on black-box cases (no problem-specific information is provided in the prompts) have been observed. Unlike previous studies that focus on optimizing a single performance criterion, MEoH (Yao et al., 2024b) considers multiple performance metrics, including optimality and efficiency, and seeks a set of trade-off algorithms in a single run in a multi-objective evolutionary framework. A dominance-dissimilarity score is designed for effectively searching the complex algorithm space.
@@
LLM-based agent design has also gained much attention. For example, ADAS (Hu et al., 2024) proposes an automated design of agentic systems, which aims to automatically generate powerful agentic system designs by using meta agents that program new agents. They present a novel algorithm, meta agent search, which iteratively creates new agents from an archive of previous designs, demonstrating through experiments that these agents can outperform state-of-the-art hand-designed agents across various domains. Further studies on LLM-based agent systems are discussed in (Sun et al., 2024a) and (Liu et al., 2024f).
@@
We identify two primary challenges in existing works on LLMaD: 1) LLMs struggle with complex algorithm design tasks. Most existing studies focus on developing specific algorithmic components, such as key heuristics and code snippets, rather than complete algorithms (Liu et al., 2024e). 2) Algorithm design is domain-specific, and LLMs typically possess limited knowledge about specific algorithm design tasks. Consequently, standalone LLMs often fall short in terms of performance (Zhang et al., 2024b). The effectiveness of LLMaD depends on search frameworks that can iteratively interact with both LLMs and their environments. Details on the search methods are provided in the next section.